NUCLEAR THEORY, Vol. 29 eds. A. Georgieva, N. Minkov, Heron Press, Sofia

Preparation of Papers for Heron Press Science Series Books

\underline{F.A. Author}^{1}, S.B. Author Jr.^{2}, C. Author^{1,2}

1Academic Affiliation, Zip Code City, State
2Academic Affiliation, Zip Code City, State

Abstract. These instructions give you guidelines for preparing papers for Heron Press Science Series Books (HPSS). Use this document as a template if you are using L^{A}TEX. Otherwise, use this document as an instruction set. Carefully follow the paper submission process in Sec. II of this document. Keep in mind that the electronic file you submit will be formatted further at Heron Press Publishing House. This first paragraph is formatted in the abstract style. Be sure to define all symbols used in the abstract, and do not cite references in this section.

Nomenclature (if needed)

et al.= et al.; use the predefined command \etal i.e.= i.e.; use the predefined command \ie e.g.= e.g.; use the predefined command \eg etc.= etc.; use the predefined command \etc vs.= vs. (versus); use the predefined command \vs A = amplitude of oscillation a = cylinder diameter

1 Introduction

Authors using L^{A}TEXwill need to save the HeronPressDoc.tex, HERON.cls, and bk10.clo files in an appropriate directory of their hard drive. To create your formatted manuscript, type your own text over sections of HeronPress-Doc.tex, or cut and paste from another document and then use the available markup styles.

2 Detailed Formatting Instructions

The styles and formats for the HPSS Template have been incorporated into the structure of this document. Please use this template to prepare your manuscript; you can simply type your own text over sections of this document, or cut and paste from another document and use the available markup styles.

F.A. Author, S.B. Author Jr., C. Author

2.1 Document Text

The default font for the HPSS Template is Times New Roman, 10-point size. There should be an additional (blank) line between paragraphs.

2.2 Headings

Format the title of your paper with capital and lower-case letters.

Use the predefined command \author{1}{2} for the responsible author, where ‘1’ stands for the name of the author and ‘2’ points to the affiliation line, and the command \coauthor{1}{2} for the other authors.

To express more information related to a given author, please use the com-

mand \thanks{ }.

To proceed with the Authors’ Index, please fill in the field of the supplied

command appropriately \index{NAME, INITIALS}.

The affiliation line (\address{1}{2} includes author’s city, state, and zip/postal code (as appropriate) and ends with a respective number.

For headings and subheadings: use commands \section, \subsection, and \subsubsection if necessary for cross-referencing sections within the

paper, and \section*, \subsection*, and \subsubsection* when

cross-referencing is not required.

2.3 Abstract

The abstract is one paragraph long (not an introduction) and complete in itself (no reference numbers). It should indicate subjects dealt with in the paper and state the objectives of the investigation. Newly observed facts and conclusions of the experiment or argument discussed in the paper must be stated in summary form; readers should not have to read the paper to understand the abstract.

2.4 References

List and number all references at the end of the paper by employing the environ-

ment \begin{thebibliography}{99}. . . \end{thebibliography}

with the help of the provided package cite. The abbreviated journal name is preferred. Use the complete date for daily and weekly publications. Corresponding numbers are used to cite references in the text [1], unless the citation is an integral part of the sentence (e.g., “It is shown in Ref. [2] that… ”) or follows a mathematical expression: “A^{2} + B = C (Ref. [3]).” For multiple citations, separate reference numbers with commas [4, 7]. Reference citations in the text should be in numerical order.

In the reference list, give all authors’ names; do not use “et al.” unless there are four authors or more. Papers that have not been published should be cited as “unpublished”; papers that have been submitted or accepted for publication should be cited as “submitted for publication.” Private communications and personal Web sites should appear as footnotes rather than in the reference list.

Preparation of Papers for Heron Press Science Series Books

References should be cited according to the standard publication reference style (for examples, see the “Refferences” section of this template). As a rule, all words are capitalized except for articles, conjunctions, and prepositions of four letters or fewer. Names and locations of publishers should be listed; month and year should be included for reports and papers. For papers published in translation journals, please give the English citation first, followed by the original foreign language citation.

2.5 Figures, Tables, and Other Images

Insert tables and figures within your document either scattered throughout the text or all together at the end of the file. Figures should have no background, borders, or outlines. Make sure the figurename is “Figure”.

Place figure captions below all figures. If your figure has multiple parts, include the labels “a),” “b),” etc.., below and to the left of each part, above the figure caption. Please verify that the figures and tables you mention in the text actually exist. When citing a figure or table in the text, do not abbreviate, use the full labels “Figure”, “Table.” Number each different type of illustration (ı.e., figures, tables, images) sequentially with relation to other illustrations of the same type.

Figure axis labels are often a source of confusion. Use words rather than symbols wherever possible. As in the Figure 1 example in this document, write the quantity “Magnetization” rather than just “M.” Do not enclose units in parentheses, but rather separate them from the preceding text by commas. Do not label axes only with units. As in Figure 1, for example, write “Magnetization, A/m” or “Magnetization, A m^{−1},” not just “A/m.” Do not label axes with a ratio of quantities and units. For example, write “Temperature, K,” not “Temperature/K.”

Multipliers can be especially confusing. Write “Magnetization, kA/m” or

Figure 1. Magnetization as a function of applied fields.

“Magnetization, 10^{3} A/m.” Do not write “Magnetization (A/m)×1000” because the reader would not then know whether the top axis label in Figure 1 meant 16000 A/m or 0.016 A/m.

See the Table 1 example for table style and column alignment. Note that table contents should be typed in \small, fill the width of the page, and the data rows should not be separate with lines.

Table 1. Transitions selected for thermometry

Transition

Line ν  J  Frequency, cm^{−1}  F J, cm^{−1}  Gν, cm^{−1}
a  0  P_{12}  2.5  44069.416  73.58  948.66
b  1  R_{2}  2.5  42229.348  73.41  2824.76
c  2  R_{21}  805  40562.179  71.37  4672.68
d  0  R_{2}  23.5  42516.527  1045.85  948.76

2.6 Equations, Numbers, Symbols, and Abbreviations

Equations are centered and numbered consecutively following the L^{A}TEXstandard procedure for typing in display mode as in Eq. (1). Do not insert a blank line on either side of the equation except a new paragraph is beginning.

T_{ij} = ρu_{i}u_{j} − λx_{i}x_{j} − ξθ(u_{i}u_{j} − g_{ij}) ,  (1)

Be sure that the symbols in your equation are defined before the equation appears, or immediately following. Italicize symbols (T might refer to temperature, but T is the unit tesla). Refer to “(1),” not “Eq. (1),” or “equation (1)” except at the beginning of a sentence: “Equation (1) is… ”.

Define abbreviations and acronyms the first time they are used in the main text. Very common abbreviations such as SI, ac, and dc do not have to be defined. Abbreviations that incorporate periods should not have spaces: write “P.R.,” not “P. R.” Delete periods between initials if the abbreviation has three or more initials; e.g., U.N. but USA. Do not use abbreviations in the title unless they are unavoidable.

3 Conclusion

Although a conclusion may review the main points of the paper, it must not replicate the abstract. A conclusion might elaborate on the importance of the work or suggest applications and extensions. Do not cite references in the conclusion as all points should have been made in the body of the paper. Note that the conclusion section is the last section of the paper to be numbered. The appendix (if present), acknowledgment, and references are listed without numbers.

Preparation of Papers for Heron Press Science Series Books

Appendix

An appendix, if needed, appears before the acknowledgements.

Acknowledgements

Avoid expressions such as “One of us (S.B.A.) would like to thank… ” Instead, write “F.A. Author thanks… ” Sponsor and financial support acknowledgments are also to be listed in the “acknowledgments” section.

References

[1] A.V. Mikhailov, Physica D 3 (1986) 1872-1873.

[2] M.A. Dornheim, Aviation Week and Space Technology 145, No. 24, 9 Dec. (1996) 44-46.

[3] F. Calogero and A. Degasperis, Lett. Nuovo Cimento 16 (1976) 425-433.

[4] V.S. Gerdjikov, G.G. Grahovski, and N.A. Kostov, Eur. Phys.J. B29 (2002) 243-248.

[5] R. Peyret and T.D. Taylor, Computational Methods in Fluid Flow, 2^{nd} ed., Springer-Verlag, New York (1983) Chaps. 7, 14.

[6] P. Ring and P. Schuk, The Nuclear Many-Body Problem, Springer, Berlin (1980).

[7] Y. Chi (ed.), Fluid Mechanics Proceedings, SP-255, NASA, 1993.

[8] V. Gerdjikov and M. Tsvetkov (eds.), Prof. G. Manev’s Legacy in Contemporary Astronomy, Theoretical and Gravitational Physics, Heron Press, Sofia (2005) 235-238.

[9] J.D. Morris, In: Proceedings of the Annual Heat Transfer Conference, editted by B. Corbell, Vol. 1, Inst. of Mechanical Engineering, New York (1992) 227-234.

[10] G.T. Chapman and M. Tobak, “Nonlinear Problems in Flight Dynamics,” NASA TM-85940 (1984).

[11] J.L. Steger, Jr., C.J. Nietubicz, and J.E. Heavey, “A General Curvilinear Grid Generation Program for Projectile Configurations,” U.S. Army Ballistic Research Lab., Rept. ARBRL-MR03142, Aberdeen Proving Ground, MD, Oct. (1981).

[12] K. Tseng, “Nonlinear Green’s Function Method for Transonic Potential Flow,” Ph.D. Dissertation, Aeronautics and Astronautics Dept., Boston Univ., Cambridge, MA (1983).

[13] E. Budevski, I. Radev, and E. Slavcheva, “New Trends in Gas Electrode Structures: the Solar Hydrogen Energy Conversion Cycles,” Suppl. Bulg. J. Phys. [CD-ROM] Vol. 32 No 4, Heron Press, Sofia (2006).

[14] C.P. Atkins and J.D. Scantelbury, “The Activity Coefficient of Sodium Chloride in a Simulated Pore Solution Environment,” Journal of Corrosion Science and Engineering [online journal], Vol. 1, No. 1, Paper 2, URL: http://www.cp/umist.

ac.uk/JCSE/vol1/vol1.html [cited 13 April 1998].

[15] A. Vickers, “10-110 mm/hr Hypodermic Gravity Design A,” Rainfall Simulation

Database [online database], URL: http://www.geog.le.ac.uk/bgrg/

lab.htm [cited 15 March 1998].

[16] TAPP, Thermochemical and Physical Properties, Software Package, Ver. 1.0, E. S. Microware, Hamilton, OH (1992).

[17] R. Scherrer, D. Overholster, and K. Watson, Lockheed Corp., Burbank, CA, U.S. Patent Application for a “Vehicle,” Docket No. P-01-1532, filed 11 Feb. (1979).

[18] J. Doe, “Title of Paper,” Name of Journal (to be published).

[19] J. Doe, “Title of Chapter,” Name of Book, edited by… Publisher’s name and location (to be published).

[20] J. Doe, “Title of Work,” Name of Archive, Univ. (or organization) Name, City, State, Year (unpublished).